A companys accounting records included the following accounts counts. Receivable (A) 434 322 2,755 a 382 12/31 30 12/31 12/31 Required: 1. For each T-account, describe the typical transactions that cause it to increase and decrease. 2. Compute the missing amounts Complete this question by entering your answers in the tabs below. Required 1 Required 2 For each T-account, describe the typical transactions that cause it to increase and decrease. Transactions Increase and Decrease Accounts Receivable Sales to customers on account Prepaid Rert Prepaid Rent Cash collections from customers Cash payments for ront related to future poriods The bonefits of these prepayments are used up ovor tme Deferred RevenueCaah recoived from customers for goods or services to tbe provided in the futute Deferred Revenue Goods and services are provided Required 2
Answer #1

1)

Accounts Transactions Increase and decrease
Account receivable Sales to customers on account Increase
Account receivable Cash collections from customers Decrease
Prepaid rent Cash payments for rent related to future periods Increase
Prepaid rent The benefits of these prepayments are used up over there Decrease
Deferred revenue Cash received from customer for goods and services to be provided in the future Increase
Deferred revenue Goods and services are provided Decrease

Calculate missing amount

a) 322+2755-382 = $2695

b) 45+87-30 = $102

c) 434+396-444 = $386